OpenStudy (anonymous):

Find the equation of the line tangent to f(x)=2x^2+2x=1 at (1,2). Walk me through please!

OpenStudy (ash2326):

we have to find tangent at (1,2) on the curve of f(x) to find the equation of the tangent , we need its slope slope is given by f'(x) at the point 1,2 or f'(1) f'(x)=4x+2 slope is value of f'(x) at the point (1,2) f'(1)=6 Tangent's slope is 6 and it passes through 1,2 we know equation of a line passing through a point x1, y1 and having slope m is given by (y-y1)/(x-x1)=m (y-2)/(x-1)=6 y-2=6x-6 or y=6x-4 this is the tangent's equation